Determine whether the function f(x) = x^2 + 3x + 2 from Z to Z is a bijective

function.


1
Expert's answer
2021-08-10T13:38:42-0400

The function "f(x)=x^2+3x+2" is not one-to-one.

For example

"0\in Z, -3\in Z"


"f(0)=(0)^2+3(0)+2=2"

"f(-3)=(-3)^2+3(-3)+2=2"

Hence


"f(0)=2=f(3)"

Since the function "f(x)=x^2+3x+2" is not one-to-one, then the given function is not bijective.